OpenStudy (anonymous):

In a classroom, 2/3 of the students have brown hair. furthermore, 2/5 have both brown hair and brown eyes. the probability that a brown haired student also has brown eyes equals ____. A. 1/2 B. 2/3 C. 1/3 D. 3/5 I keep getting a different answer than the answers there. What am I doing wrong?

OpenStudy (ybarrap):

Bayes Theorem: P(brown eyes|brown hair) = P(brown eyes & brown hair)/P(brown hair) =(2/5)/(2/3) =3/5 Make sense?

OpenStudy (anonymous):

yes thank you! I doing it backwards

OpenStudy (ybarrap):

yw